# Feasibility of Learning Weighted Automata on a Semiring

### v3

@@ 3.2. Guessable and strongly guessable functions
The definition of weakly guessable functions ensures the existence of some finite sets $Q,T$ and coefficients $\lambda\in\Lambda_{Q,T}$ that together determine a correct hypothesis automaton. In other words, functions that are not weakly guessable are those for which all hypothesis automata are incorrect, and so (in fairness to the learner!) one should immediately exclude all such functions from consideration if one seeks to learn by membership and equivalence in the style of Angluin. For weakly guessable functions, the problem then becomes: how to find a suitable $Q,T$ and $\lambda$. Clearly $Q$ must be a row-generating set for the full Hankel matrix; but is it always possible to find such a set from finitely many membership and equivalence queries? Also, even if a suitable row generating set $Q$ is found, there is no guarantee that accompanying $T$ and $\lambda$ can be found in finite time. Guessable functions are those for which there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. Strongly guessable functions are those for which for every finite row-generating set $Q$ there exists a finite subset $T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. In the following definition, the reader should recall from Remark 2.6 that the sets $\Lambda_{Q,T}$ depend upon the function $f$ (and hence also the semiring $S$).
@@
A function $f:\Sigma^{*}\rightarrow S$ is guessable if there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. We say that the pair $(Q,T)$ witnesses that $f$ is guessable. Furthermore, we say that a guessable function $f$ is strongly guessable if for all finite sets $Q\subseteq\Sigma^{*}$ with $\Lambda_{Q}\neq\emptyset$ there exists a finite set $T\subseteq\Sigma^{*}$ such that $(Q,T)$ witnesses that $f$ is guessable.
@@
Let $Q$ be a finite subset of $\Sigma^{*}$. A total $Q$-chain is an infinite descending chain of solution spaces $\Lambda_{Q,T_{0}}\supsetneq\Lambda_{Q,T_{1}}\supsetneq\Lambda_{Q,T_{2}}\supsetneq\cdots$ with $T_{i}\subseteq T_{i+1}$ for all $i\geq 0$ and such that each word in $\Sigma^{*}$ is contained in some $T_{i}$. We say that a function $f:\Sigma^{*}\rightarrow S$ is row-bound if there exists a finite set of words $Q$ such that $\Lambda_{Q}\neq\emptyset$ and there is no total $Q-$chain. We say that $f:\Sigma^{*}\rightarrow S$ is strongly row-bound if for every finite set of words $Q$ such that $\Lambda_{Q}\neq\emptyset$ there is no total $Q$-chain.
The above definition is to be linked with the progress measure defined in [21, Definition 11]. The existence of a progress measure for a Hankel matrix implies that the corresponding function is strongly row-bound. However, the converse might not be true. A progress measure implies a uniform bound, for all $Q$, on the number of elements in a total $Q$-chain, but for a (strongly) row-bound function, even if no such chain can be infinite, its size might depend on the set of words $Q$. The progress measure is a stronger assumption, and one of the sufficient condition for termination of a learning algorithm [21], while the property of being (strongly) row-bounded pinpoints more precisely a characterisation of (strongly) guessable functions. We give a variant of the notion of progress measure in the next subsection.
@@
Let $f:\Sigma^{*}\to S$ be a weakly guessable function. Then $f$ is (strongly) guessable if and only if $f$ is (strongly) row-bound.
@@
Suppose that $Q,T$ are finite sets of words such that $\Lambda_{Q,T}=\Lambda_{Q}$. Let $(T_{i})_{i\geq 0}$ be a sequence of subsets of $\Sigma^{*}$ with $T_{i}\subseteq T_{i+1}$ for all $i$ and such that $\bigcup_{i}T_{i}=\Sigma^{*}$. By definition, there exists some set $T_{j}$ such that $T\subseteq T_{j}$. Then for all $i\geq j$ we have $T\subseteq T_{j}\subseteq T_{i}$. Since $\Lambda_{Q,T}=\Lambda_{Q}$ it then follows that $\Lambda_{Q,T_{i}}=\Lambda_{Q,T_{j}}=\Lambda_{Q,T}$ for all $i\geq j$. Thus there can be no total $Q$-chains. It now follows that if $f$ is guessable then $f$ is row-bound, whilst if $f$ is strongly guessable, then it must be strongly row-bound.
Suppose now that $f$ is (strongly) row-bound. Let $Q$ be a finite set of words such that $\Lambda_{Q}\neq\emptyset$ and there is no total $Q$-chain. Then taking $T_{i}$ to be the finite set of words of length at most $i$ for each positive integer $i$, we have that $T_{i}\subseteq T_{i+1}$ and $\bigcup_{i}T_{i}=\Sigma^{*}$ and so there must exist $n$ such that for all $m\geq n$, $\Lambda_{Q,T_{m}}=\Lambda_{Q,T_{n}}$. Since the $T_{i}$ enumerate all words, we must therefore also have $\Lambda_{Q,T_{n}}=\Lambda_{Q}$. Hence for any $\lambda$ in $\Lambda_{Q,T_{n}}$, the hypothesis automaton over $Q,T_{n}$ and $\lambda$ computes $f$. Thus, if $f$ is (strongly) row-bound, we conclude that $f$ is (strongly) guessable. ∎
@@
As in Remarks 2.11 and 3.7, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions of co-guessable, strongly co-guessable: see Appendix A for full details. With these definitions, one can prove the following entirely analogous result.
@@
Let $f:\Sigma^{*}\to S$ be a weakly co-guessable function. Then $f$ is (strongly) co-guessable if and only if $f$ is (strongly) column-bound.
@@
Let $S$ be a field and $f:\Sigma^{*}\rightarrow S$ a function computed by a finite automaton with weights from $S$. By Remark 2.13 there exist finite sets of words $Q,P\subseteq\Sigma^{*}$ such that $Q$ is row-generating for $f$ and $P$ is column-generating for $f$. It is straightforward to check that for any such sets $Q$ and $P$ we have $\Lambda_{Q,P}=\Lambda_{Q}\neq\emptyset$ (see Lemma 4.1 below for a generalisation of this fact) and hence $f$ is strongly guessable (and by a dual argument, $f$ is strongly co-guessable).
3.3. Learning algorithms
As mentioned earlier, if one knows that the target function is weakly guessable, that is to say that there is a hypothesis automaton that computes it, the question becomes to find some suitable $Q,T$ and $\lambda$. If for each pair $Q,T$ the set $\Lambda_{Q,T}$ is finite (as is the case for any finite semiring), then by enumerating the set of pairs $(Q,T)$, one can construct a naive learning algorithm for weakly guessable functions given in Figure LABEL:fig:algoweak; we say that such functions have finite solvability.
Similarly, by enumerating the set of pairs $(Q,T)$, one can construct (regardless of the cardinalities of the sets $\Lambda_{Q,T}$) the naive learning algorithm for guessable functions given in Figure LABEL:fig:algoguess.
For strongly guessable functions, Figure LABEL:fig:algostrong gives an alternative algorithm in which $Q$ and $T$ are enumerated and incremented alternately.
The applicability of each of these algorithms is dependent on decidability of equivalence queries (although, in practice, it may be possible to approximate equivalence queries using many membership queries) and solvability of general finite $S$-linear systems of equations. Indeed, a requirement for the algorithm to function is to be able to determine whether left-linear combinations exist, and when this is the case, to be able to find at least one such. This is given in the solvability condition.
@@ 3.3. Learning algorithms
A function $f$ is solvable if there is an algorithm that, for any finite sets $Q$, $T$, can determine whether $\Lambda_{Q,T}\neq\emptyset$ and compute at least one element of $\Lambda_{Q,T}$ in that case.
Clearly, if $S$ is a semiring where there exists an algorithm which can take as input an arbitrary system of linear equations over $S$ and compute a solution, then all functions over the semiring are solvable. We note that in [21] such semirings were called solvable. It is clear that all finite semirings are solvable (a very naive algorithm would be to try all possibilities) and fields where the basic operations are computable are also solvable (via Gaussian elimination). In [21] the authors provide a sufficient condition for a principal ideal domain to be solvable. Turning to other examples, we note for example that over $\mathbb{Z}_{\rm max}$ there is a straightforward (and well-known) method to determine whether a system of the form $A\otimes x=b$ has a solution. Indeed, if $x$ is to be a solution, we must in particular have that $A_{i,j}+x_{j}\leq b_{i}$ for all $i,j$. If $A_{i,j}=-\infty$, then this inequality holds automatically, otherwise, we require that $x_{j}\leq b_{i}-A_{i,j}$. In order to solve the inequality, for each $j$ such that there exists at least one $i$ with $A_{i,j}\neq-\infty$ we could set $x_{j}={\rm min}\{b_{i}-A_{i,j}:A_{i,j}\neq-\infty\}$, and set all remaining $x_{j}$ equal to 00. It is clear that this is a computable (we use finitely many operations of subtraction and comparison on finitely many entries drawn from $\mathbb{Z}_{\rm max}$) solution to the inequality $A\otimes x\leq b$. In some sense it can be viewed as a “greatest” solution since the entries $x_{j}$ where there exists $A_{i,j}\neq-\infty$ cannot be taken any larger (of course, if there are any columns of $A$ where the entries are all equal to $\infty$- ∞, the corresponding entry of $x$ is irrelevant and our choice of taking $x_{j}=0$ was completely arbitrary). Furthermore, it can also be seen that either the solution to the inequality is a solution to the equation $A\otimes x=b$, or else no solution to the equation $A\otimes x=b$ exists. (Indeed, if the solution to the inequality we constructed does not satisfy the equation, then we must have that for some $j$ the strict inequality holds: ${\rm max}(A_{i,j}+x_{j})<b_{i}$, meaning that either all $A_{i,j}$ are equal to $\infty$- ∞ and so there is no hope to satisfy the equation ${\rm max}(A_{i,j}+x_{j})<b_{i}$, or else in order to attain $A_{i,j}+x_{j}=b_{i}$ we require that one of the $x_{j}$ with $A_{i,j}$ must take a larger value, which as we have seen, is not possible.) Thus $\mathbb{Z}_{\rm max}$ (and likewise $\mathbb{N}_{\rm max}$) is solvable.
We also note that each of the naive algorithms outlined above do not make any use of the data provided by the counterexamples. We do not develop further algorithms here: as mentioned in the introduction, our main aim in this paper is to picture the landscape for active learning in arbitrary semirings, and demonstrate the limitations. Our intention is to lay the foundations for future progress, indicating clear avenues for future investigation. However, to conclude this section we reproduce, in Figure LABEL:fig:algo, the general algorithm given in [21], and discuss the relationship to the present work.
In [21], an example showing that this algorithm does not always terminate is given. This example corresponds to the automaton given in Figure 1(d), that (as we shall see in the next section) is not weakly guessable. Since the algorithm can only guess hypothesis automata, it is clear that the algorithm will not terminate on any function that is not weakly guessable. Moreover, if the function is not guessable, then (by definition) for each finite set of words $Q$ with $\Lambda_{Q}\neq\emptyset$ one has that $\Lambda_{Q}\neq\Lambda_{Q,T}$ for all finite sets of words $T$. This in turn means that there is no guarantee that the algorithm will terminate, since it could be that in every test step the chosen $\lambda$ generates an incorrect hypothesis automaton.
In [21], three conditions are requested to ensure termination and correctness of the algorithm in Figure LABEL:fig:algo: the ascending chain condition, the progress measure and the solvability condition. Below we give slight variants on these conditions, that ensure that the proof given in  [21] still holds (and any function satisfying these conditions will be guessable). The following strong ascending chain condition ensures that the closure step always terminates in finite time for a fixed $T$ and that it cannot be taken infinitely many times for different $T$’s.
@@
A function $f:\Sigma^{*}\to S$ with Hankel matrix $F$ satisfies the strong ascending chain condition if for all sequences of left-semimodules $(X_{i})_{i\in\mathbb{N}}$ generated by finite sets of rows of $F$, such that $X_{0}\subseteq X_{1}\subseteq X_{2}\subseteq\cdots$ there is $n$ such that for all $m\geq n$, $X_{m}=X_{n}$.
Meanwhile the following progress measure ensures that one cannot go through the Test step infinitely many times without changing the set $Q$ in the closure step.
Definition 3.18.
Let $f$ be a function computed by a weighted automaton over $S$, let :=$\mathcal{L}:=\{(Q,T):Q,T\mbox{ are finite sets of words and $Q$ is left-closedon $T$}\}$. We say that $f$ has a progress measure if there is a totally ordered set $(\mathbb{K},\succ)$ with no infinite decreasing chain and a function $\mu:\mathcal{L}\rightarrow\mathbb{K}$ with the property that whenever $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$ we have $\mu(Q,T)\succ\mu(Q,T^{\prime})$.
It is clear from the definitions that if $f$ has a progress measure then it must be strongly row-bound. Moreover, with these definitions, the following proposition now derives in a similar manner to results presented in [21].

### v4

@@ 3.2. Guessable and strongly guessable functions
The definition of weakly guessable functions ensures the existence of some finite sets $Q,T\subseteq\Sigma^{*}$ and coefficients $\underline{\lambda}\in\Lambda_{Q,T}$ (writing here and from now on simply $Q$ in place of the set of rows of the Hankel matrix indexed by the words in $Q$, as discussed in Notation 2.6) that together determine a correct hypothesis automaton. In other words, functions that are not weakly guessable are those for which all hypothesis automata are incorrect, and so (in fairness to the learner!) one should immediately exclude all such functions from consideration if one seeks to learn by membership and equivalence in the style of Angluin. For weakly guessable functions, the problem then becomes: how to find a suitable $Q,T$ and $\lambda$. Clearly $Q$ must be a row-generating set for the full Hankel matrix; but is it always possible to find such a set from finitely many membership and equivalence queries? Also, even if a suitable row generating set $Q$ is found, there is no guarantee that accompanying $T$ and $\lambda$ can be found in finite time. Guessable functions are those for which there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. Strongly guessable functions are those for which for every finite row-generating set $Q$ there exists a finite subset $T\subseteq\Sigma^{*}$ such that all the associated hypothesis automata compute the correct function. In the following definition, the reader should recall from Remark 2.7 that the sets $\Lambda_{Q,T}$ depend upon the function $f$ (and hence also the semiring $S$).
@@
A function $f:\Sigma^{*}\rightarrow S$ is guessable if there exist finite subsets $Q,T\subseteq\Sigma^{*}$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. We say that the pair $(Q,T)$ witnesses that $f$ is guessable. Furthermore, we say that a guessable function $f$ is strongly guessable if for all finite sets $Q\subseteq\Sigma^{*}$, there exists a finite set $T\subseteq\Sigma^{*}$ such that $\Lambda_{Q,T}=\Lambda_{Q}$.
@@
Let $Q$ be a finite subset of $\Sigma^{*}$. A total $Q$-chain is an infinite descending chain of left-solution sets $\Lambda_{Q,T_{0}}\supsetneq\Lambda_{Q,T_{1}}\supsetneq\Lambda_{Q,T_{2}}\supsetneq\cdots$ with $T_{i}\subseteq T_{i+1}$ for all $i\geq 0$ and such that each word in $\Sigma^{*}$ is contained in some $T_{i}$. We say that a function $f:\Sigma^{*}\rightarrow S$ is row-bound if there exists a finite set of words $Q$ such that $\Lambda_{Q}\neq\emptyset$ and there is no total $Q-$chain. We say that $f:\Sigma^{*}\rightarrow S$ is strongly row-bound if for every finite set of words $Q$, there is no total $Q$-chain.
The above definition is related to the progress measure defined in [36, Definition 11]; we will discuss this in the next subsection.
@@
Let $f:\Sigma^{*}\to S$. Then $f$ is guessable if and only if $f$ is row-bound. Moreover, $f$ is strongly guessable if and only if $f$ is strongly row-bound and weakly guessable.
@@
Suppose that $Q,T$ are finite sets of words such that $\Lambda_{Q,T}=\Lambda_{Q}$. Let $(T_{i})_{i\geq 0}$ be a sequence of subsets of $\Sigma^{*}$ with $T_{i}\subseteq T_{i+1}$ for all $i$ and such that $\bigcup_{i}T_{i}=\Sigma^{*}$. By definition, there exists some set $T_{j}$ such that $T\subseteq T_{j}$. Then for all $i\geq j$ we have $T\subseteq T_{j}\subseteq T_{i}$. Since $\Lambda_{Q,T}=\Lambda_{Q}$ it then follows that $\Lambda_{Q,T_{i}}=\Lambda_{Q,T_{j}}=\Lambda_{Q,T}$ for all $i\geq j$. Thus there can be no total $Q$-chains. It now follows that if $f$ is guessable then $f$ is row-bound, whilst if $f$ is strongly guessable, then it must be strongly row-bound (and weakly guessable).
Now, let $Q$ be a finite set of words such that there is no total $Q$-chain. Then taking $T_{i}$ to be the finite set of words of length at most $i$ for each positive integer $i$, we have that $T_{i}\subseteq T_{i+1}$ and $\bigcup_{i}T_{i}=\Sigma^{*}$ and so there must exist $n$ such that for all $m\geq n$, $\Lambda_{Q,T_{m}}=\Lambda_{Q,T_{n}}$. Since the $T_{i}$ enumerate all words, we must therefore also have $\Lambda_{Q,T_{n}}=\Lambda_{Q}$. Thus if $f$ is row-bound, there is such a $Q$ such that $\Lambda_{Q}\neq\emptyset$ and we conclude that $f$ is guessable. Additionally, if $f$ is strongly row-bound and weakly guessable then the property holds for all $Q$ and at least one of them is such that $\Lambda_{Q}\neq\emptyset$, which makes $f$ strongly guessable.
@@
As in Remarks 2.14 and 3.8, one can also work with the transpose of the Hankel matrix, leading to obvious left-right dual notions of co-guessable, strongly co-guessable: see Appendix A for full details. With these definitions, one can prove the following entirely analogous result.
@@
Let $f:\Sigma^{*}\to S$. Then $f$ is (strongly) co-guessable if and only if $f$ is (strongly) column-bound and weakly co-guessable.
3.3. Algorithmic strategies
Game strategies.
We now explain what the notions of (weakly/strongly) guessable mean in terms of strategies for the learner to discover an automaton correctly computing the target function via a sequence of Angluin-style queries.
The hierarchy we propose is (in general) orthogonal to any complexity classes - in particular there is no guarantee that a strongly guessable function can be learnt in polynomial time. See Remark 3.20 at the end of the section.
@@ 3.3. Algorithmic strategies
Provided the semiring S is countable, a trivial algorithm would be to enumerate all the automata and make an equivalence query on each of them in turn. One would of course like to improve upon this exponential time strategy. Even though there is no guarantee in terms of time complexity of the Angluin-style process we describe next, note that there are two main differences: (1) the learning process still applies when the semiring is uncountable, (2) it will only explore, in a prescribed manner, a subset of automata i.e. the class of hypothesis automata.
Consider a two player game between the teacher and the learner, in which the learner’s objective is to find an automaton computing a target function. The rules are as follows. At every round, the learner is allowed to:
  • choose a pair of finite subsets $Q,T\subseteq\Sigma^{*}$;
  • ask the teacher to provide an hypothesis automaton constructed from $Q$ and $T$.
The teacher must:
  • declare if there are no possible hypothesis automaton for a given pair $Q,T$ (that is, if $\Lambda_{Q,T}=\emptyset$) or provide a hypothesis automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ for this pair (by selecting $\underline{\lambda}\in\Lambda_{Q,T}$);
  • declare whether each provided automaton computes the target function and if not, provide a counter-example.
The game stops when the teacher declares an automaton that computes the target function. We view the teacher as an oracle, able to determine an element of each of the sets $\Lambda_{Q,T}$, $\Lambda_{Q}$ and $\Lambda_{Q,T}\setminus\Lambda_{Q}$ (for all finite subsets of words $Q,T$), or else determine that the set is empty. In this way, all computation is deferred to the teacher. The teacher must always answer truthfully, but may play as an ally (meaning whenever $(Q,T)$ is queried with $\Lambda_{Q}\neq\emptyset$, the teacher will choose to reply with a correct automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ where $\underline{\lambda}\in\Lambda_{Q}$) or as an adversary (whenever $(Q,T)$ is queried with $\Lambda_{Q,T}\neq\Lambda_{Q}$, the teacher will choose to reply with an incorrect automaton $\mathcal{H}_{Q,T,\underline{\lambda}}$ where $\underline{\lambda}\in\Lambda_{Q,T}\setminus\Lambda_{Q}$).
One can record the history of a game up to any given point as the finite (possibly empty) sequence of pairs $(Q,T)$ queried so far, together with the replies from the teacher (i.e. the hypothesis automata and counter-examples or the fact that $\Lambda_{Q,T}=\emptyset$). A strategy for the learner is a function mapping each history $h$ where the last pair in the sequence does not signify the end of the game to a new pair $(Q_{h},T_{h})\in\Sigma^{*}\times\Sigma^{*}$. Additionally, the strategy is said to be incremental if for all histories $h$, $Q_{h}\supseteq Q$ and $T_{h}\supseteq T$ holds for all pairs $(Q,T)$ already queried in $h$. A strategy is winning for the learner, if, for all possible plays following that strategy, the teacher eventually declares that a correct automaton is found.
Proposition 3.17.
A function is weakly guessable if and only if the learner has a winning strategy in the game where the teacher plays as an ally.
Proof.
For a weakly guessable function there exists a non-empty finite subset $Q\subseteq\Sigma^{*}$ such that $\Lambda_{Q}\neq\emptyset$. Since $\Sigma^{*}$ is countable, so is the set of non-empty finite subsets of $\Sigma^{*}$. Let $(Q_{i})_{i\in\mathbb{N}}$ be an enumeration of the finite non-empty subsets of $\Sigma^{*}$ and consider the following strategy for the learner: if $h$ is empty, query $(Q_{0},\{\varepsilon\})$; if the last query of $h$ is $(Q_{i},\{\varepsilon\})$ and does not halt the game then query $(Q_{i+1},\{\varepsilon\})$ (no matter the automata and counter-examples given by the teacher). The learner will eventually query $(Q,\{\varepsilon\})$ and since $\Lambda_{Q}\subseteq\Lambda_{Q,\{\varepsilon\}}$ the teacher (playing as an ally) will choose to declare a correct hypothesis automaton. Conversely, if the learner has a winning strategy, then clearly there exist finite sets $Q,T\subseteq\Sigma^{*}$ and $\underline{\lambda}\in\Lambda_{Q,T}$ such that $\mathcal{H}_{Q,T,\underline{\lambda}}$ computes the target function and so by definition the function is weakly guessable. ∎
Proposition 3.18.
A function is guessable if and only if the learner has a winning strategy in the game where the teacher plays as an adversary.
Proof.
For a guessable function there exists finite subsets $Q,T\subseteq\Sigma^{*}$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. Since $\Sigma^{*}$ is countable, so is the set of pairs of non-empty finite subsets of $\Sigma^{*}$; let $(Q_{i},T_{i})_{i\in\mathbb{N}}$ be an enumeration of all such pairs and consider the following strategy for the learner: if the history is empty, query $(Q_{0},T_{0})$; if the last query of $h$ is $(Q_{i},T_{i})$ and does not halt the game then query $(Q_{i+1},T_{i+1})$ (no matter the automata and counter-examples given by the teacher). The learner will eventually query $(Q,T)$ and since $\Lambda_{Q}=\Lambda_{Q,T}$, the teacher will be forced to declare a correct hypothesis automaton.
Conversely, if the learner has a winning strategy in the game where the teacher plays as an adversary, then clearly there exist finite sets $Q,T\subseteq\Sigma^{*}$ and $\underline{\lambda}\in\Lambda_{Q,T}$ such that $\mathcal{H}_{Q,T,\underline{\lambda}}$ computes the target function and $\Lambda_{Q,T}\setminus\Lambda_{Q}=\emptyset$ (otherwise the teacher can always choose to provide an incorrect automaton). So by definition the function is guessable. ∎
Proposition 3.19.
If a function is strongly guessable, then the learner has a winning strategy that is incremental in the game where the teacher plays as an adversary.
Proof.
Let $Q_{i}=T_{i}$ be the set of all words of length at most $i$, and consider the following strategy for the learner: if the history is empty, query $(Q_{0},T_{0})$; if $(Q_{i},T_{j})$ is the last pair queried in $h$ and the teacher declared that $\Lambda_{Q_{i},T_{j}}$ is non-empty, but did not give a correct automaton, then query $(Q_{i},T_{j+1})$; if $(Q_{i},T_{j})$ is the last pair queried in $h$ and $\Lambda_{Q_{i},T_{j}}$ is empty then query $(Q_{i+1},T_{j})$; otherwise if $(Q_{i},T_{j})$ is the last pair queried in $h$ and the teacher declared a correct automaton, the game halts.
For any given target function this process is completely deterministic, but possibly infinite: since the teacher plays as an adversary (and so never supplies a correct automaton if there is an incorrect choice), the sequence of pairs queried by the learner for this function will be the same regardless of different choices of automata or counterexample supplied by the teacher in response to each query (only the question of their existence is relevant to proceedings). Our aim is to show that whenever the target function is strongly guessable, this process terminates.
Since $f$ is weakly guessable and for all non negative integers $i$, $Q_{i}\subseteq Q_{i+1}$ then there exists a non negative integer $I$ such that for all $i<I$, $\Lambda_{Q_{i}}$ is empty and $\Lambda_{Q_{I}}$ is non-empty. Hence by strong guessability, for all $i<I$, there exists $T$ such that $\Lambda_{Q_{i},T}$ is empty and there exists $T^{\prime}$ such that $\Lambda_{Q_{I},T^{\prime}}=\Lambda_{Q_{I}}$ is non-empty. By Remark 2.8 and Lemma 2.9, it follows that for all $i<I$, there exists $\ell$ such that for all $j\geq\ell$, $\Lambda_{Q_{i},T_{j}}$ is empty and $J$ such that for all $j\geq J$, $\Lambda_{Q_{I},T_{j}}=\Lambda_{Q_{I}}$ is non-empty (and none of the $\Lambda_{Q_{I},T}$ is empty for any $T$). Hence, the process terminates, and the startegy above is winning for the learner. ∎
Remark 3.20.
The strategies mentioned above do not guarantee any polynomial time complexity. For strongly guessable functions, a polynomial time strategy would require the learner to make more well-informed choices of $Q$ and $T$ (by making use of the counter-examples) so that the size of $Q$ and $T$ remain polynomial with respect to the size of the automaton to learn and counter-examples given. For guessable and weakly guessable functions, in addition to the constraints on the size of $Q$ and $T$, the learner needs to be able to find a correct $\underline{\lambda}$ in polynomial time; this may be possible over certain semirings.
Relationship to the algorithmic approach of [36].
In [36], a general learning algorithm is given (subject to the existence of a ‘closedness strategy’) and three conditions (solvability, ascending chain condition, and progress measure) are requested to ensure termination and correctness of the algorithm for all functions computed by weighted automata over a fixed semiring. We give an overview of this algorithm using our terminology in Figure LABEL:fig:algo, and discuss below some weaker conditions that are easily seen to be sufficient for the proof of termination and correctness on a given target function, and how these relate to our definitions of guessability.
First, note that this algorithm requires access to a procedure (referred to as a ‘closedness strategy’ in [36]) that can (a) determine whether $Q$ is left-closed on $T$ (or in other words, whether $\Lambda_{Q,T}\neq\emptyset$); (b) if $\Lambda_{Q,T}=\emptyset$, can find $q\in Q$ and $a\in\Sigma$ such that the corresponding equation has no solution; and (c) if $\Lambda_{Q,T}\neq\emptyset$, can find an element $\underline{\lambda}$ of $\Lambda_{Q,T}$.
Clearly, if $S$ is a semiring where there exists an algorithm which can take as input an arbitrary system of linear equations over $S$ and compute a solution, then there is a closedness strategy for all functions over this semiring. We note that in [36] such semirings were called solvable. It is clear that all finite semirings are solvable (a very naive algorithm would be to try all possibilities) and fields where the basic operations are computable are also solvable (via Gaussian elimination). In [36] the authors provide a sufficient condition for a principal ideal domain to be solvable (specifically: all ring operations are computable, and each element can be effectively factorised into irreducibles).
Remark 3.21.
Turning to other examples, we note over $\mathbb{Z}_{\rm max}$ there is a straightforward (and well-known) method to determine whether a system of the form $A\otimes\underline{x}=\underline{b}$ has a solution. Indeed, if $\underline{x}$ is a solution, we must in particular have that $A_{i,j}+\underline{x}_{j}\leq\underline{b}_{i}$ for all $i,j$. If $A_{i,j}=-\infty$, then this inequality holds automatically, otherwise, we require that $\underline{x}_{j}\leq\underline{b}_{i}-A_{i,j}$. In order to solve the inequality, for each $j$ such that there exists at least one $i$ with $A_{i,j}\neq-\infty$ we could set $\underline{x}_{j}={\rm min}\{\underline{b}_{i}-A_{i,j}:A_{i,j}\neq-\infty\}$, and set all remaining $\underline{x}_{j}$ equal to 00. It is clear that this is a computable (we use finitely many operations of subtraction and comparison on finitely many entries drawn from $\mathbb{Z}_{\rm max}$) solution to the system of inequalities $A\otimes\underline{x}\leq\underline{b}$. Furthermore, it can also be seen that either this solution is also a solution to the equation $A\otimes\underline{x}=\underline{b}$, or else no solution to the equation $A\otimes\underline{x}=\underline{b}$ exists. (Indeed, if the solution to the inequality that we have constructed does not satisfy the equation, then we must have that for some $i$ the following strict inequality holds: ${\rm max}(A_{i,j}+\underline{x}_{j})<\underline{b}_{i}$, meaning that either all $A_{i,j}$ are equal to $\infty$- ∞ and so there is no hope to satisfy the corresponding equation, or else in order to attain the equality we require that one of the $\underline{x}_{j}$ for $j$ with $A_{i,j}\neq-\infty$ must take a larger value than set above, which by construction is not possible.) Thus $\mathbb{Z}_{\rm max}$ (and likewise $\mathbb{N}_{\rm max}$) is solvable.
Remark 3.22.
In [36], an example showing that the algorithm in Figure LABEL:fig:algo does not always terminate is given. This example corresponds to the automaton given in Figure 1(d), that (as we shall see in the next section) is not weakly guessable. Since the algorithm can only guess hypothesis automata, it is clear that the algorithm will not terminate on any function that is not weakly guessable. In fact, termination cannot be ensured if the target function is not guessable. Indeed, if the function is not guessable, then (by definition) for each finite set of words $Q$ with $\Lambda_{Q}\neq\emptyset$ one has that $\Lambda_{Q}\neq\Lambda_{Q,T}$ for all finite sets of words $T$. This in turn means that there is no guarantee that the algorithm will terminate, since it could be that in every test step the element $\underline{\lambda}$ provided by the closedness strategy generates an incorrect hypothesis automaton.
For the remainder of this section, let us assume that $f$ is a function for which a closure strategy exists. The following condition ensures that the closure step terminates in finite time for each fixed $T$ and that it cannot be taken infinitely many times for different $T$’s.
@@
A function $f:\Sigma^{*}\to S$ with Hankel matrix $F$ satisfies the ascending chain condition on rows if for all sequences of left-semimodules $(X_{i})_{i\in\mathbb{N}}$ generated by finite sets of rows of $F$, such that $X_{0}\subseteq X_{1}\subseteq X_{2}\subseteq\cdots$ there is $n$ such that for all $m\geq n$, $X_{m}=X_{n}$.
Finally, we note that the following condition (which mimics the definition of progress measure given in [36], rewritten in the language of the present paper and extracting the key property) ensures that one cannot go through the Test step infinitely many times without changing the set $Q$ in the closure step.
@@
Let $f$ be a function computed by a weighted automaton over $S$ and let
:=$\mathcal{L}:=\{(Q,T):Q,T\mbox{ are finite sets of words and }Q\mbox{ is left-closed on }T\}.$
We say that $f$ has a weak progress measure if there is a totally ordered set $(\mathbb{K},\prec)$ with no infinite descending chain and a function $\mu:\mathcal{L}\rightarrow\mathbb{K}$ with the property that if $T\subseteq T^{\prime}$ and $\Lambda_{Q,T}\supsetneq\Lambda_{Q,T^{\prime}}$ we have $\mu(Q,T)\succ\mu(Q,T^{\prime})$.
We summarise the observations above in the following proposition, which provides a slight generalisation of statement of the main result of [36]; the proof derives in a similar manner to results presented in [36].